### v1

# On the Friendly number of 10

Abstract.
To answer the question ”Is $10$ a solitary number?”, we have established a necessary and sufficient condition on primes $p$ and $q$ so that $p$ divides $\sigma(q^{2a})$($a\in\mathbb{Z^{+}}$). Also We have found the form of $q$ whenever $p$ divides $\sigma(q^{2a})$ for $q>p$ Using these two, we have proved that if $n$ is a friend of $10$ then it is a odd square with at least seven distinct prime factors, $5$ being the least one. Moreover there exists a prime factor $p$ of $n$ such that $2a+1\equiv 0\pmod{f}$ and $5^{f}\equiv 1\pmod{p}$ where $f$ is the smallest odd positive integer greater than $1$ and less than or equal to $\min\{2a+1,p-1\}$, provided $5^{2a}\mid\mid n$, $a\in\mathbb{Z^{+}}$. Further, there exist prime factors $p$ and $q$ not necessarily distinct of $n$ such that $p\equiv 1\pmod{10}$ and $q\equiv 1\pmod{6}$. Besides, we have proved that if a Fermat prime $F_{k}$ divides $n$ then $n$ must have a prime factor congruent to $1$ modulo $2F_{k}$.
footnotetext: Mathematics Subject Classification (2020):11A25.footnotetext: Keywords: Abundancy Index, Sum of Divisors,Friendly Numbers, Solitary Numbers.
1. Introduction
For any positive integer $n$, $\sigma(n)$ denotes the sum of the positive divisors of $n$. The abundancy index of $n$ is defined as $\frac{\sigma(n)}{n}$. We use $I(n)=\frac{\sigma(n)}{n}$ to denote the abundancy index throughout in this paper. A positive integer $n$ is said to be friendly number if there exist a positive integer $m$ other than $n$ such that it has same abundancy index as $n$ i.e., $I(n)=I(m)$ and in that case we say $m$ is a friend of $n$ and if $n$ has no friend then $n$ is called a solitary number. In [2] we can find that if $(\sigma(n),n)=1$ then $n$ is a solitary number. For a prime number $p$ we have $\sigma(p)=p+1$ thus $(\sigma(p),p)=1$ which shows that any prime $p$ is a solitary number. Further using this we can show that all powers of primes are in fact solitary. $10$ is the smallest positive integer whose classification in terms of solitary and friendly is unknown as in case of $1,2,3,4,5,7,8,9$ we can use $(\sigma(n),n)=1$ this fact to prove them solitary and for $6$ it is well known that it has many friends $28,496,8128$ etc. (all friends of $6$ are called perfect numbers). Although all friends of $6$ are found even and therefore one question naturally raised that ”are there any odd perfect numbers?” which is still an open problem in mathematics. However in this paper we will discuss on the friend of $10$ mainly and we will prove the following theorems.
Theorem 1.1.
[…]
Corollary 1.2.
Let $p$, $p^{*}$ and $q$ be three distinct prime numbers with $p^{k-1}\mid\mid(q-1)$ and $p^{*k^{*}-1}\mid\mid(q-1)$ $(k,k^{*}\in\mathbb{Z^{+}}$) also let $p\mid\sigma(q^{2a})$ and $p^{*}\mid\sigma(q^{2a^{*}})$( $a^{*}$$\in\mathbb{Z^{+}}$) with $f_{p}^{q}$, $f_{p^{*}}^{q}$ respectively in Theorem 1.1. If $f_{p^{*}}^{q}\mid f_{p}^{q}$ then $pp^{*}\mid\sigma(q^{2a})$.
Corollary 1.3.
Let $n$ be a friend of $10$ with $5^{2a}\mid\mid n$, $a\in\mathbb{Z^{+}}$. Then there exists a prime factor $p$ of $n$ such that $2a+1\equiv 0\pmod{f}$ and $5^{f}\equiv 1\pmod{p}$ where $f$ is the smallest odd positive integer greater than $1$ and less than or equal to $\min\{2a+1,p-1\}$.
Corollary 1.4.
Let $n$ be a friend of $10$. If any Fermat prime $F_{k}=2^{2^{k}}+1$ ($k\in\mathbb{Z}_{\geq 0}$) divides $n$ then there exists a prime factor $p$ of $n$ such that $p\equiv 1\pmod{2F_{k}}$.
Theorem 1.5.
If $p$ and $q$ are two distinct prime numbers with $q>p$, then $p$ divides $\sigma(q^{2a}),\leavevmode\nobreak\ a\in\mathbb{Z^{+}}$ if and only if for $r\neq 1$, $q\equiv r\pmod{p}$ and $2a+1\equiv 0\pmod{f}$ where $f$ is the smallest odd positive integer greater than $1$ such that $r^{f}\equiv 1\pmod{p}$ and for $r=1$, $q\equiv r\pmod{p}$ and $2a+1\equiv 0\pmod{p}$.
Corollary 1.6.
If $n$ is a friend of $10$ then $n$ has a prime factor $p$ such that $p\equiv 1\pmod{6}$.
Theorem 1.7.
If $n$ is a friend of $10$, then $n$ is a odd square with at least $7$ distinct prime factors, and $5$ is the smallest among them. Further, there exists prime factor $p$ and $q$ not necessarily distinct of $n$ such that $p\equiv 1\pmod{10}$ and $q\equiv 1\pmod{6}$.
Before jumping into the proofs, let’s first discuss some preliminaries.
2. Properties of the Abundancy Index
In this section, we will prove some elementary properties of the abundancy index, which will not only help us to understand the behavior of the abundancy index but also help us to prove some properties of the friend of $10$. These properties will play a crucial role in proving our main theorems.
Lemma 2.1.
$I(n)$ is weakly multiplicative.
Proof.
Let $a$ and $b$ be two positive integers such that $(a,b)=1$. Then $\sigma(ab)=\sigma(a)\sigma(b)$ as $\sigma(n)$ is a weakly multiplicative function and therefore $I(ab)=\frac{\sigma(ab)}{ab}=\frac{\sigma(a)\sigma(b)}{ab}=\frac{\sigma(a)}{a}\frac{\sigma(b)}{b}=I(a)I(b)$. This proves that $I(n)$ is weakly multiplicative. ∎
Lemma 2.2.
Let $a,n$ be two positive integers and $a>1$. Then $I(an)>I(n)$.
Proof.
Let $a,n$ be two positive integers with $a>1$. Then
$\sigma(an)$ $=\sum_{d\mid an}d=\sum_{d^{{}^{\prime}}\mid n}ad^{{}^{\prime}}+\sum_{{\begin{subarray}{c}d^{{}^{\prime\prime}}\mid an\\ d^{{}^{\prime\prime}}\neq ad^{{}^{\prime}}\\ \text{where\leavevmode\nobreak\ }d^{{}^{\prime}}\mid n\end{subarray}}}d^{{}^{\prime\prime}}$
$=a\sum_{d^{{}^{\prime}}\mid n}d^{{}^{\prime}}+\sum_{{\begin{subarray}{c}d^{{}^{\prime\prime}}\mid an\\ d^{{}^{\prime\prime}}\neq ad^{{}^{\prime}}\\ \text{where\leavevmode\nobreak\ }d^{{}^{\prime}}\mid n\end{subarray}}}d^{{}^{\prime\prime}}$
$=a\sigma(n)+\sum_{{\begin{subarray}{c}d^{{}^{\prime\prime}}\mid an\\ d^{{}^{\prime\prime}}\neq ad^{{}^{\prime}}\\ \text{where\leavevmode\nobreak\ }d^{{}^{\prime}}\mid n\end{subarray}}}d^{{}^{\prime\prime}}.$
Therefore $\frac{\sigma(an)}{a\sigma(n)}>1$ and this proves $I(an)>I(n)$. ∎
Lemma 2.3.
Let $p_{1}$, $p_{2}$, $p_{3}$,…, $p_{m}$ be $m$ distinct primes and $a_{1}$, $a_{2}$, $a_{3}$ ,…, $a_{m}$ are positive integers then
$I\biggl{(}\prod_{i=1}^{m}p_{i}^{a_{i}}\biggl{)}=\prod_{i=1}^{m}\biggl{(}\sum_{j=0}^{a_{i}}p_{i}^{-j}\biggl{)}=\prod_{i=1}^{m}\frac{p_{i}^{a_{i}+1}-1}{p_{i}^{a_{i}}(p_{i}-1)}.$
Proof.
Since
$\sigma\biggl{(}\prod_{i=1}^{m}p_{i}^{a_{i}}\biggl{)}=\prod_{i=1}^{m}\sigma(p_{i}^{a_{i}})=\prod_{i=1}^{m}\biggl{(}\sum_{j=0}^{a_{i}}p_{i}^{j}\biggl{)}=\prod_{i=1}^{m}\frac{p_{i}^{a_{i}+1}-1}{p_{i}-1}$
We have
$I\biggl{(}\prod_{i=1}^{m}p_{i}^{a_{i}}\biggl{)}=\frac{\sigma\biggl{(}\prod_{i=1}^{m}p_{i}^{a_{i}}\biggl{)}}{\prod_{i=1}^{m}p_{i}^{a_{i}}}=\prod_{i=1}^{m}\frac{p_{i}^{a_{i}+1}-1}{p_{i}^{a_{i}}(p_{i}-1)}=\prod_{i=1}^{m}\biggl{(}\sum_{j=0}^{a_{i}}p_{i}^{-j}\biggl{)}$
This completes the proof. ∎
Lemma 2.4.
If $p_{1}$,…,$p_{m}$ are distinct primes, $q_{1}$,…,$q_{m}$ are distinct primes such that $p_{i}\leq q_{i}$ for all $1\leq i\leq m$. If $t_{1},t_{2},...,t_{m}$ are positive integers then
$I\biggl{(}\prod_{i=1}^{m}p_{i}^{t_{i}}\biggl{)}\geq I\biggl{(}\prod_{i=1}^{m}q_{i}^{t_{i}}\biggl{)}.$
Proof.
For $1\leq i\leq m$, let us define $f_{i}:(1,\infty)\longrightarrow\mathbb{R}$ by $f_{i}(x)=\frac{x^{t_{i}+1}-1}{x^{t_{i}+1}-x^{t_{i}}}$, for some positive integer $t_{i}$. Then
$f_{i}^{{}^{\prime}}(x)=\frac{t_{i}x^{t_{i}-1}(x-1)-x^{t_{i}}(x^{t_{i}}-1)}{(x^{t_{i}+1}-x^{t_{i}})^{2}}$
$=\frac{x^{t_{i}-1}(x-1)\{t_{i}-x(x^{t_{i}-1}+x^{t_{i}-2}+...+1)\}}{(x^{t_{i}+1}-x^{t_{i}})^{2}}.$
Clearly, as $x>1,t_{i}\geq 1$, $t_{i}x^{t_{i}-1}(x-1)>0$ and $(x^{t_{i}+1}-x^{t_{i}})^{2}>0$ but $t_{i}-x(x^{t_{i}-1}+x^{t_{i}-2}+...+1)\}<0$, therefore we have $f_{i}^{{}^{\prime}}(x)<0$ for all $x\in(1,\infty)$. This shows that $f_{i}$ is a decreasing function for each $x\in(1,\infty)$ and for all $1\leq i\leq m$. Now let $p_{i}$, $q_{i}\in(1,\infty)$ such that $p_{i}\leq q_{i}$, for all $1\leq i\leq m$. Then for all $1\leq i\leq m$, $f_{i}(p_{i})\geq f_{i}(q_{i})$ i.e.,
$\frac{p_{i}^{t_{i}+1}-1}{p_{i}^{t_{i}}(p_{i}-1)}\geq\frac{q_{i}^{t_{i}+1}-1}{q_{i}^{t_{i}}(p_{i}-1)}$
Therefore
$\prod_{i=1}^{m}\frac{p_{i}^{t_{i}+1}-1}{p_{i}^{t_{i}}(p_{i}-1)}\geq\prod_{i=1}^{m}\frac{q_{i}^{t_{i}+1}-1}{q_{i}^{t_{i}}(p_{i}-1)}$
implies
$\frac{\sigma\biggl{(}\prod_{i=1}^{m}p_{i}^{t_{i}}\biggl{)}}{\prod_{i=1}^{m}p_{i}^{t_{i}}}\geq\frac{\sigma\biggl{(}\prod_{i=1}^{m}q_{i}^{t_{i}}\biggl{)}}{\prod_{i=1}^{m}q_{i}^{t_{i}}}.$
This proves $I\biggl{(}\prod_{i=1}^{m}p_{i}^{t_{i}}\biggl{)}\geq I\biggl{(}\prod_{i=1}^{m}q_{i}^{t_{i}}\biggl{)}$.
Lemma 2.5.
If $n=\prod_{i=1}^{m}p_{i}^{a_{i}}$, then $I(n)<\prod_{i=1}^{m}\frac{p_{i}}{p_{i}-1}$.
Proof.
Since, for every $1\leq i\leq m$ we have
$\frac{p_{i}^{a_{i}+1}-1}{p_{i}^{a_{i}}(p_{i}-1)}<\frac{p_{i}^{a_{i}+1}}{p_{i}^{a_{i}}(p_{i}-1)}=\frac{p_{i}}{p_{i}-1}$
Therefore
$\prod_{i=1}^{m}\frac{p_{i}^{a_{i}+1}-1}{p_{i}^{a_{i}}(p_{i}-1)}<\prod_{i=1}^{m}\frac{p_{i}}{p_{i}-1}$
i.e., $I(n)<\prod_{i=1}^{m}\frac{p_{i}}{p_{i}-1}$ and this proofs the lemma.
3. Basic Properties of friend of 10
In this section, we will prove some basic properties of the friend of $10$, which will help us to understand the nature of the friend of $10$. Moreover these results will play a significant role in the proofs of our main theorems.
Proposition 3.1.
Let $n$ be a friend of $10$ then $n$ is an odd positive square integer.
Proof.
Let a positive integer $n$ be a friend of 10 then $n$ must be greater than $10$ as for any positive integer less than $10$, $I(n)\neq\frac{9}{5}$. Since $I(n)=\frac{9}{5}$ i.e., $\frac{\sigma(n)}{n}=\frac{9}{5}$ i.e., $5\sigma(n)=9n$, we have $5\mid n$ and $9\mid\sigma(n)$ as $(5,9)=1$, therefore we can write $n=5n^{\prime}$ for some positive integer $n^{\prime}>2$. If possible, let $n^{\prime}$ be an even positive integer. Then we can rewrite $n$ as $n=10n^{\prime\prime}$ for some positive integer $n^{\prime\prime}>1$, but then $I(n)>I(10)$ by Lemma 2.2. Therefore we can conclude that $(n^{\prime},2)=1$ i.e., $n$ is an odd positive integer. Now if $n=5^{a}$ for some positive integer $a>2$ then it can not be a friend of $10$ as by Lemma 2.5, $I(5^{a})<\frac{5}{4}$. Therefore $n$ must be written as $n=5^{a_{1}}\prod_{i=2}^{m}p_{i}^{a_{i}}$($p_{1}=5$) where $p_{i}$ are prime numbers and $a_{i}$ are positive integers. If possible, suppose that some of $a_{i}$ are not even (without loss of generality let $a_{r}$ be odd, $1\leq r\leq m$ ). Then
$I(n)=\frac{\sigma(n)}{n}=\frac{9}{5}$
By Lemma 2.1, we have
$I(5^{a_{1}})\cdot\prod_{i=2}^{m}I(p_{i}^{a_{i}})=\frac{9}{5}$
This implies
$\sigma(5^{a_{1}})\cdot\prod_{i=2}^{m}\sigma(p_{i}^{a_{i}})=9\cdot 5^{{a_{1}}-1}\cdot\prod_{i=2}^{m}p_{i}^{a_{i}},$
that is we get
$(1+5+...+5^{a_{1}})\cdot\cdot\cdot\cdot(1+p_{r}+...+p_{r}^{a_{r}})\cdot\cdot\cdot(1+p_{m}+...+p_{m}^{a_{m}})$
$=9\cdot 5^{{a_{1}}-1}\cdot\prod_{i=2}^{m}p_{i}^{a_{i}}.$
Since $p_{i}>2$, for all $1\leq i\leq m$ the right hand side of the above expression is an odd integer but the left hand side is even since $a_{r}$ is odd which immediately implies $(1+p_{r}+...+p_{r}^{a_{r}})$ is even. Hence all $a_{i}$ must be even integers. This proves that $n$ is an odd positive square integer. ∎
Proposition 3.2.
Let $n$ be a friend of $10$, then $5$ is the smallest prime factor of it.
Proof.
Let $n$ be a friend of $10$, then by Proposition 3.1, $n$ is an odd square therefore $3$ is the only odd prime number less than $5$. Therefore we need to check whether $3$ divides $n$ or not. Suppose that $3$ divides $n$. Then we can write $n$ as $n=5^{2a}\cdot 3^{2b}\cdot m^{2}$ where $a,b,m\in\mathbb{Z^{+}}$ with $(m,5)=(m,3)=1$. Therefore we have $I(5^{2a}\cdot 3^{2b}\cdot m^{2})\geq I(5^{2a}\cdot 3^{2b})$ by Lemma 2.2. Now if any of $a,b$ is greater than $1$ then $I(5^{2a}\cdot 3^{2b})>\frac{9}{5}$ i.e., $I(n)>\frac{9}{5}$, therefore the only possibility is $a=b=1$. But if $a=b=1$ then
$I(5^{2}\cdot 3^{2}\cdot m^{2})=\frac{9}{5}$
By Lemma 2.1, we have
$I(5^{2})\cdot I(3^{2})\cdot I(m^{2})=\frac{9}{5}$
i.e.,
$31\cdot 13\cdot\sigma(m^{2})=9\cdot 5\cdot 3^{2}\cdot m^{2}$
This shows that $31\cdot 13\mid m^{2}$, therefore rewriting $n$ as $n=5^{2}\cdot 3^{2}\cdot 13^{2c}\cdot 31^{2d}\cdot t^{2}$ where $c$, $d$, $t\in\mathbb{Z^{+}}$. But again by Lemma 2.2 we have that $I(n)=I(5^{2}\cdot 3^{2}\cdot 13^{2c}\cdot 31^{2d}\cdot t^{2})\geq I(5^{2}\cdot 3^{2}\cdot 13^{2}\cdot 31^{2})=\frac{20191}{10075}>\frac{9}{5}$. Therefore we can conclude that $3$ doesn’t divide $n$. This proves that $5$ is the least prime factor of $n$. ∎
4. Proofs of the Main Theorems
4.1. Proof of Theorem 1.1.
Let $p$ and $q$ be two distinct primes with $p^{k-1}\mid\mid(q-1)$ where $k$ is some positive integer. Suppose that $p$ divides $\sigma(q^{2a})$, $a\in\mathbb{Z^{+}}$. Then
$\sigma(q^{2a})\equiv 0\pmod{p}$
i.e.,
$\sum_{r=0}^{2a}q^{r}\equiv 0\pmod{p}$
i.e.,
$\frac{q^{2a+1}-1}{q-1}\equiv 0\pmod{p}$
Therefore,
$q^{2a+1}\equiv 1\pmod{p^{k}}$
By $Euler-Fermat$ theorem we can write
[…]
Let us assume that $f$ be the smallest positive integer such that $q^{f}\equiv 1\pmod{p^{k}}$ then $f>1$ as if $f=1$ then $q\equiv 1\pmod{p^{k}}$ i.e., $p^{k}\mid(q-1)$ which is impossible as $p^{k-1}\mid\mid(q-1)$. Since $2a+1\geq f$ we can write $2a+1=tf+r$ for some $t\in\mathbb{Z^{+}}$ and $0\leq r<f$. Then
$1\equiv q^{2a+1}\equiv q^{tf+r}\equiv q^{r}\pmod{p^{k}}$
[…]
Conversely, let $2a+1\equiv 0\pmod{f}$ and $q^{f}\equiv 1\pmod{p^{k}}$ where $f$ is the smallest odd positive integer greater than $1$. Then $2a+1=tf$, for some $t\in\mathbb{Z^{+}}$ and therefore
$q^{2a+1}=q^{tf}\equiv 1\pmod{p^{k}}$
Since $(q-1,p^{k})=p^{k-1}$, we have
$\frac{q^{2a+1}-1}{(q-1)}\equiv 0\pmod{p}$
This implies
$\sigma(q^{2a})\equiv 0\pmod{p}$
In other words, $p$ divides $\sigma(q^{2a})$ and this proves the theorem.∎
Remark 4.1.
If $n=\prod_{i=1}^{m}p_{i}^{2a_{i}}$$(p_{1}=5)$ is a friend of 10 then $I(n)=\frac{9}{5}$ i.e., $\sigma(5^{2a_{1}})\cdot\prod_{i=2}^{m}\sigma(p_{i}^{2a_{i}})=9\cdot 5^{{2a_{1}}-1}\cdot\prod_{i=2}^{m}p_{i}^{2a_{i}}$. This shows that for every prime factors $p_{i}$ of $n$, there exists some $p_{j}$ other than $p_{i}$ such that $p_{i}\mid\sigma(p_{j}^{2a_{j}})$ and also for every $p_{i}$, the prime factors of $\sigma(p_{i}^{2a_{i}})$ are precisely from the set $\{3,5,p_{2},p_{3},...,p_{m}\}.$
4.2. Proof of Corollary 1.2.
Let $p$, $p^{*}$ and $q$ be three distinct primes and $p\mid\sigma(q^{2a})$ and $p^{*}\mid\sigma(q^{2a^{*}})$ with $f_{p}^{q}$, $f_{p^{*}}^{q}$ respectively in Theorem 1.1. Also suppose that $f_{p^{*}}^{q}\mid f_{p}^{q}$. Then by Theorem 1.1 we have $2a+1\equiv 0\pmod{f_{p}^{q}}$ and $2a^{*}+1\equiv 0\pmod{f_{p^{*}}^{q}}$ and since $f_{p^{*}}^{q}\mid f_{p}^{q}$ we can write $2a+1\equiv 0\pmod{f_{p^{*}}^{q}}$. This shows that $p^{*}\mid\sigma(q^{2a})$ and since $(p,p^{*})=1$, we have $pp^{*}\mid\sigma(q^{2a})$.∎
4.3. Proof of Corollary 1.3.
Let $n=\prod_{i=1}^{m}p_{i}^{2a_{i}}$$(p_{1}=5)$ with $5^{2a}\mid\mid n$, ($a\in\mathbb{Z^{+}}$) be a friend of $10$. In particular let $q=5$ in Theorem 1.1. Then there must exists a prime $p$ of $n$ such that $p\mid\sigma(5^{2a})$. Since $p\geq 5$, $p\nmid 4$, therefore we have $2a+1\equiv 0\pmod{f}$ and $5^{f}\equiv 1\pmod{p}$ where $f$ is the smallest odd positive integer greater than 1. As $f\mid(2a+1)$ and $f\mid\phi(p)$ i.e., $f\mid(p-1)$, we have $f\leq\min\{2a+1,p-1\}.$
4.4. Proof of Corollary 1.4.
Let $n$ be a friend of $10$ and $F_{k}=2^{2^{k}}+1$ ($k\in\mathbb{Z}_{\geq 0}$) be a prime factor of $n$. If possible, let $n$ has no prime factor congruent to $1$ modulo $F_{k}$. Now since $F_{k}\mid n$, by Theorem 1.1 there exists a prime factor $p$ of $n$ with $p^{2a_{p}}\mid\mid n$ ($a_{p}\in\mathbb{Z^{+}}$) and $F_{k}^{k^{*}-1}\mid\mid(p-1)$ such that $2a_{p}+1\equiv 0\pmod{f}$ and $p^{f}\equiv 1\pmod{F_{k}^{k^{*}}}$ where $f$ is smallest positive odd integer greater than $1$. Since no prime factors $p$ of $n$ is congruent to $1$ modulo $F_{k}$, we have $k^{*}=1$. Further as $f\mid\phi(F_{k}^{k^{*}})$ i.e., $f\mid\phi(F_{k})$ and $\phi(F_{k})=2^{2^{k}}$ this shows that there does not exist any odd positive divisor of $\phi(F_{k})$ greater than $1$ as it has only one prime factor $2$. Therefore by Theorem 1.1 we can conclude that $F_{k}$ does not divide any $\sigma(p^{2a_{p}})$. Since $n$ is a friend of 10 using Remark 4.1 we can conclude that $F_{k}$ can not be a divisor of $n$ but this contradicts our assumption that $F_{k}$ is a prime factor of $n$. Hence there must exists a prime factor $p$ of $n$ such that $p\equiv 1\pmod{F_{k}}$. Since $(p-1)$ is even we have $p\equiv 1\pmod{2F_{k}}$.∎
Remark 4.2.
If $n$ is a friendly number of $10$ then it has a prime factor $p$ such that $p\equiv 1\pmod{10}$ as $5\mid n$ and $5$ is a Fermat prime. Also it is useful to note that as $17$ is a Fermat prime and if $17\mid n$ then $n$ must have a prime factor $p$, such that $p\equiv 1\pmod{34}$.
4.5. Proof of Theorem 1.5
Let $p$ and $q$ be two distinct prime numbers ($q>p$) such that $p\mid\sigma(q^{2a})$ for some positive integer $a$. Then
$\sigma(q^{2a})\equiv 0\pmod{p}$
i.e.,
(4.1) $\sum_{j=0}^{2a}q^{j}\equiv 0\pmod{p}$
As $q>p$, we can write $q=pk+r$, for some $k\in\mathbb{Z^{+}}$ and $0\leq r\leq(p-1)$. If $r=0$ then $q=pk$ but $p\nmid\sigma((pk)^{2a})$, therefore $1\leq r\leq(p-1)$. From (4.1) we have that
$\sum_{j=0}^{2a}(pk+r)^{j}\equiv 0\pmod{p}$
i.e.,
$\sum_{j=0}^{2a}r^{j}\equiv 0\pmod{p}$
If $r=1$ then
$2a+1\equiv 0\pmod{p}$
And if $r\neq 1$ then
$\frac{r^{2a+1}-1}{r-1}\equiv 0\pmod{p}$
Since $(r-1,p)=1$ we have
$r^{2a+1}\equiv 1\pmod{p}$
By $Euler-Fermat$ theorem we can write
$r^{p-1}\equiv 1\pmod{p}$
Let $f$ be the smallest positive integer such that $r^{f}\equiv 1\pmod{p}$. Then $f>1$ otherwise $r\equiv 1\pmod{p}$ i.e., $r=1$. Since $2a+1\geq f$ we can write $2a+1=tf+s$ for some $t\in\mathbb{Z^{+}}$ and $0\leq s<f$. Then
$1\equiv r^{2a+1}\equiv r^{tf+s}\equiv r^{s}\pmod{p}$
[…]
Conversely, let for $r\neq 1$, $q\equiv r\pmod{p}$ and $2a+1\equiv 0\pmod{f}$ where $f$ is the smallest odd positive integer greater than 1 such that $r^{f}\equiv 1\pmod{p}$ and for $r=1$, $q\equiv r\pmod{p}$ and $2a+1\equiv 0\pmod{p}$.
For $r\neq 1$, since $r^{f}\equiv 1\pmod{p}$, we have $q^{f}\equiv r^{f}\equiv 1\pmod{p}$. If possible there exists $f^{\prime}<f$ such that $q^{f^{\prime}}\equiv 1\pmod{p}$ then $1\equiv q^{f^{\prime}}\equiv r^{f^{\prime}}\pmod{p}$. This shows that $f^{\prime}>f$ by our assumption on $f$. But this leads to a contradiction to that fact that $f^{\prime}<f$, therefore $f$ is the smallest odd positive integer such that $q^{f}\equiv 1\pmod{p}$ with $2a+1\equiv 0\pmod{f}$. Since $p\nmid(q-1)$, by Theorem 1.1 we can conclude that $p\mid\sigma(q^{2a})$.
For $r=1$, we have $q\equiv 1\pmod{p}$ and $2a+1\equiv 0\pmod{p}$. For any positive integer $j$ we have
$q^{j}\equiv 1\pmod{p}$
Therefore,
$1+\sum_{j=1}^{2a}q^{j}\equiv 1+\sum_{j=1}^{2a}1\pmod{p}$
i.e.,
$\sum_{j=0}^{2a}q^{j}\equiv 2a+1\pmod{p}$
i.e.,
$\sigma(q^{2a})\equiv 0\pmod{p}$
In other words, $p$ divides $\sigma(q^{2a})$ and this completes the proof.∎
4.6. Proof of Corollary 1.6.
Let $n$ be a friend of $10$ then we have $9\mid\sigma(n)$ by Proposition 3.1. This shows that there exists a prime factor $p$ of $n$ with $p^{2a_{p}}\mid\mid n$ such that $3\mid\sigma(p^{2a_{p}})$. Since all prime factors of $n$ are greater than $3$, we can use Theorem 1.5. Now we will prove that $p$ can not be congruent to $2$ modulo $3$. Suppose that $p\equiv 2\pmod{3}$ then we must have an odd positive integer $f>1$ such that $2^{f}\equiv 1\pmod{3}$ and $2a_{p}+1\equiv 0\pmod{f}$. But no such $f$ exists as for any odd positive integer $f$ we have $2^{f}\equiv 1\pmod{3}$ i.e., $(-1)^{f}\equiv 1\pmod{3}$ i.e., $-1\equiv 1\pmod{3}$ which is absurd. Therefore $p$ must be congruent to $1$ modulo $3$. Since $p-1$ is even we have $p\equiv 1\pmod{6}$.∎
We will use Corollary 1.2 heavily to prove Theorem 1.7. We will direct use Corollary 1.2 in all cases. For example we will write ”if $11\mid\sigma(5^{2a})$ then $71\mid\sigma(5^{2a})$” as $f_{11}^{5}=5=f_{71}^{5}$ and $f_{11}^{5}\mid f_{71}^{5}$ in that case we have $2a+1\equiv 0\pmod{5}$, again ”if $19\mid\sigma(5^{2a})$ then $31\mid\sigma(5^{2a})$” as $f_{19}^{5}=9$, $f_{31}^{5}=3$ and $f_{31}^{5}\mid f_{19}^{5}$ in that case we have $2a+1\equiv 0\pmod{9}$. The values of $f_{p}^{q}$ which we will use in proving Theorem 1.7, can be found after the conclusion part.
4.7. Proof of Theorem 1.7.
Let $n$ be a friend of $10$. In Proposition 3.1 we have seen that $n=5^{a}$($a\in\mathbb{Z^{+}}$) can’t be a friend of $10$ and $n$ is an odd positive square integer therefore $n$ has at least two distinct prime factors.
Case-1:
Suppose that $n$ has exactly two distinct odd prime factors $p_{1}=5$ and $p_{2}$ where $p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}$ for some positive integers $a_{1}$, $a_{2}$. By Lemma 2.4 we have $I(5^{2a_{1}}\cdot 7^{2a_{2}})\geq I(5^{2a_{1}}\cdot p_{2}^{2a_{2}})=I(n)$ but Lemma 2.5 shows that $\frac{9}{5}>\frac{5}{4}\cdot\frac{7}{6}=\frac{35}{24}>I(5^{2a_{1}}\cdot 7^{2a_{2}})$ i.e., $\frac{9}{5}>I(5^{2a_{1}}\cdot p_{2}^{2a_{2}})=I(n)$. Therefore $n$ can’t be a friend of $10$ and $n$ has more than two distinct prime factors.
Case-2:
Suppose that $n$ has exactly three distinct odd prime factors $p_{1}=5$, $p_{2}$ and $p_{3}$ with $p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}$ where $a_{1}$, $a_{2}$, $a_{3}$ are positive integers. Again by Lemma 2.4 and Lemma 2.5 we have $I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}})\geq I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}})=I(n)$ and $\frac{9}{5}>\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{11}{10}=\frac{77}{48}>I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}})$ i.e., $\frac{9}{5}>I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}})=I(n)$. Therefore $n$ can’t be a friend of $10$ and $n$ has more than there distinct prime factors.
Case-3:
Suppose that $n$ has exactly four distinct odd prime factors $p_{1}=5$, $p_{2}$, $p_{3}$ and $p_{4}$ with $p_{4}>p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}$ where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$ are positive integers. Again by Lemma 2.4 and Lemma 2.5 we have $I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}})\geq I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}})=I(n)$ and $\frac{9}{5}>\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{11}{10}\cdot\frac{13}{12}=\frac{1001}{576}>I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}})$ i.e., $\frac{9}{5}>I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}})=I(n)$. Therefore $n$ can’t be a friend of $10$ and $n$ has more than four distinct prime factors.
Case-4:
Suppose that $n$ has exactly five distinct odd prime factors $p_{1}=5$, $p_{2}$, $p_{3}$, $p_{4}$ and $p_{5}$ with $p_{5}>p_{4}>p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}}$ where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$ are positive integers. Now we will see what are the possible values of $p_{i}$ for $2\leq i\leq 5$. If possible, let $p_{2}>7$ but then by Lemma 2.4 and Lemma 2.5, $I(n)=I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}})\leq I(5^{2a_{1}}\cdot 11^{2a_{1}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}})\leq I(5^{2a_{1}}\cdot 11^{2a_{1}}\cdot 13^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}})<\frac{5}{4}\cdot\frac{11}{10}\cdot\frac{13}{12}\cdot\frac{17}{16}\cdot\frac{19}{18}=\frac{46189}{27648}<\frac{9}{5}.$ Hence only possible value of $p_{2}$ is $7$. If possible, let $p_{3}>11$ then by Lemma 2.4 and Lemma 2.5, $I(n)\leq I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 13^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}})\leq\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{13}{12}\cdot\frac{17}{16}\cdot\frac{19}{18}=\frac{146965}{82944}<\frac{9}{5}$. Hence only possible value of $p_{3}$ is $11$. If possible, let $p_{4}>13$ then again by Lemma 2.4 and Lemma 2.5 $I(n)\leq I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}})<\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{11}{10}\cdot\frac{17}{16}\cdot\frac{19}{18}=\frac{24871}{13824}<\frac{9}{5}$. Hence only possible value of $p_{4}$ is $13$. Now if $p_{5}>23$ then by the similar argument we can show that $I(n)<\frac{9}{5}$. Therefore $p_{2}=7$, $p_{3}=11$, $p_{4}=13$ and $p_{5}\in\{17,19,23\}$ are necessary for $I(n)\geq\frac{9}{5}$. Now $p_{5}$ can’t be $17$ by Remark 4.2 as then no $p_{i}$ will be congruent to $1$ modulo $34$. If $p_{5}=23$ then by Corollary 1.3, $p_{3}=11$ is the only prime factor which divides $\sigma(5^{2a_{1}})$ but it allows $71\mid\sigma(5^{2a_{1}})$ and this is absurd as $p_{i}\neq 71$, for $1\leq i\leq 5$. Therefore only possible value of $p_{5}$ is $19$. In this case $p_{3}=11$ and $p_{5}=19$ can divides $\sigma(5^{2a_{1}})$. If $p_{3}=11\mid\sigma(5^{2a_{1}})$ then $71\mid\sigma(5^{2a_{1}})$ but $p_{i}\neq 71$, for $1\leq i\leq 5$. Therefore $p_{5}=19$ must divides $\sigma(5^{2a_{1}})$ but again it allows $31\mid\sigma(5^{2a_{1}})$ which is impossible as $p_{i}\neq 31$, for $1\leq i\leq 5$. Therefore, no prime factor of $n$ can divides $\sigma(5^{2a_{1}})$ that is $\sigma(5^{2a_{1}})$ has prime factors other than $3$, $p_{1}=5$, $p_{2}$, $p_{3}$, $p_{4}$, $p_{5}$. This proves that $n$ can’t be a friend of $10$ and hence $n$ has more than five distinct prime factors.
Case-5:
Suppose that $n$ has exactly six distinct odd prime factors $p_{1}=5$, $p_{2}$, $p_{3}$, $p_{4}$, $p_{5}$ and $p_{6}$ with $p_{6}>p_{5}>p_{4}>p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Applying similar argument given in Case-4 to get possible values of $p_{i}$, $2\leq i\leq 6$ we obtain the following ,
Figure 1.
Figure 2.
There are $19$ possible chain of values of $p_{i}$, $2\leq i\leq 6$ so that $I(n)=I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}\cdot p_{5}^{2a_{5}}\cdot p_{6}^{2a_{6}})<\frac{9}{5}$ does not occur for all $2a_{i}$ i.e., for other combination of $p_{i}$ we always have $I(n)<\frac{9}{5}$. We will begin with chain(1) and We will show that the abundancy index of these possible chain of $p_{i}$ even fails to give $\frac{9}{5}$.
Subcase-1: For chain(1), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 13^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}}\cdot p_{6}^{2a_{6}}$, where $23\leq p_{6}\leq 61$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. But by Remark 4.2 there must exists a prime factor of $n$ congruent to 1 modulo 34 as $17\mid n$ but, for any values of $p_{6}$, $n$ has no such prime factor. Therefore $n$ can’t be a friend of $10$.
Subcase-2: For chain(2), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 13^{2a_{3}}\cdot 17^{2a_{4}}\cdot 23^{2a_{5}}\cdot p_{6}^{2a_{6}}$, where $29\leq p_{6}\leq 37$ and $a_{1},a_{2},a_{3},a_{4},a_{5},a_{6}$ are positive integers. By the same argument given in Subcase-1, we can conclude that $n$ can’t be a friend of $10$.
Subcase-3: For chain(3), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 13^{2a_{3}}\cdot 19^{2a_{4}}\cdot 23^{2a_{5}}\cdot p_{6}^{2a_{6}}$, where $29\leq p_{6}\leq 31$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. But $p_{6}=29$ is impossible as by Remark 4.2, there must exists a prime factor of $n$ congruent to $1$ modulo $10$. Therefore $p_{6}=31$ Since $13\nmid(p_{i}-1)$ and $\phi(13)=12$, $f_{13}^{p_{i}}$ is $3$, for all $1\leq i\leq 6$. But $p_{i}^{3}\not\equiv 1\pmod{13}$, for each $p_{i}$, $1\leq i\leq 6$. Therefore by Theorem 1.1 we have that, $13$ does not divide any $\sigma(p_{i}^{2a_{i}})$ and which is absurd by Remark 4.1. Therefore $n$ can’t be a friend of $10$.
Subcase-4: For chain(4), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 17^{2a_{4}}\cdot 19^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $23\leq p_{6}\leq 2039$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Since $17\mid n$ by Remark 4.2 there must exists a prime factor of $n$ congruent to $1$ modulo $34$, therefore $p_{6}$ must be congruent to $1$ modulo $34$. Now $5$ can divides $\sigma(11^{2a_{3}})$ and $\sigma(p_{6}^{2a_{6}})$. But if , $5\mid\sigma(11^{2a_{3}})$ then $3221\mid\sigma(11^{2a_{3}})$ that is $p_{6}$ must be $3221$. But $p_{6}$ is not congruent $1$ modulo $34$. Hence 5 must divides $\sigma(p_{6}^{2a_{6}})$. By Proposition 3.1 ($9\mid\sigma(n)$), $3$ must divides one of $\sigma(p_{i}^{2a_{i}})$, for $1\leq i\leq 6$. But $3\nmid\sigma(p_{i}^{2a_{i}})$, for $i=1,3,4$. If $3\mid\sigma(19^{2a_{5}})$ then $127\mid\sigma(19^{2a_{5}})$ that is $p_{6}$ must be $127$ but again it is not congruent $1$ mod $34$. Now it is clear that $3\mid\sigma(7^{2a_{2}})$, we will show that $9$ can’t divides $\sigma(7^{2a_{2}})$. If $9\mid\sigma(7^{2a_{2}})$ then this immediately implies $37\mid\sigma(7^{2a_{2}})$ that is $p_{6}$ must be $37$ but it is not congruent $1$ modulo $34$. Therefore $\sigma(p_{6}^{2a_{6}})$ must be divisible by $3$ as $9\leavevmode\nobreak\ |\leavevmode\nobreak\ \sigma(n)$. Hence $p_{6}$ must be congruent to $1$ modulo $34$ and $\sigma(p_{6}^{2a_{6}})$ must be divisible by $3$ and $5$. Using Theorem 1.5 we get, $p_{6}$ is congruent to $1$ modulo $510$. Therefore $p_{6}$ either $1021$ or $1531$. If $p_{6}=1021$ then $5\mid\sigma(1021^{2a_{6}})$ but this immediately allows $41\mid\sigma(1021^{2a_{6}})$ Which is impossible as $p_{i}\neq 41$, for all $1\leq i\leq 6$. Hence the only choice for $p_{6}$ is $1531$ but again $5\mid\sigma(1531^{2a_{6}})$ implies $691\mid\sigma(1531^{2a_{6}})$ which is impossible by the same argument. Therefore there is no such $p_{6}$ for which $5$ can divides $\sigma(p_{6}^{2a_{6}})$. Hence $n$ can’t be a friend of $10$.
Subcase-5: For chain(5), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 17^{2a_{4}}\cdot 23^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $29\leq p_{6}\leq 97$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. By the same argument given in Subcase-1, we can conclude that $n$ can’t be a friend of $10$.
Subcase-6: For chain(6), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 17^{2a_{4}}\cdot 29^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $31\leq p_{6}\leq 47$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. By the same argument given in Subcase-1, we can conclude that $n$ can’t be a friend of $10$.
Subcase-7: For chain(7), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 17^{2a_{4}}\cdot 31^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $37\leq p_{6}\leq 43$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. By the same argument given in Subcase-1, we can conclude that $n$ can’t be a friend of $10$.
Subcase-8: For chain(8), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 19^{2a_{4}}\cdot 23^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $29\leq p_{6}\leq 59$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Now there are only three prime factors of $n$, $p_{3}=11$, $p_{4}=19$, $p_{6}=31$ or $p_{6}=59$ that can divide $\sigma(5^{2a_{1}})$. Now if $p_{3}=11\mid\sigma(5^{2a_{1}})$, immediately $71\mid\sigma(5^{2a_{1}})$ again if $p_{4}=19\mid\sigma(5^{2a_{1}})$, immediately $829\mid\sigma(5^{2a_{1}})$. Therefore $p_{3}=11$, $p_{4}=19$ can’t divide $\sigma(5^{2a_{1}})$ as for both the cases $p_{6}$ must be $71$ and $829$ respectively but $p_{6}\leq 59$. Therefore $p_{6}$ must divides $\sigma(5^{2a_{1}})$ and $p_{6}$ must be either $31$ or $59$ as for other values of $p_{6}$, no $p_{i}$ will divides $\sigma(5^{2a_{1}})$ If $p_{6}=31$ then $31\mid\sigma(5^{2a_{1}})$ allows $829\mid\sigma(5^{2a_{1}})$ but it is impossible as $p_{i}\neq 829$, for all $1\leq i\leq 6$. Therefore only possible value of $p_{6}$ is $59$. But $35671\mid\sigma(5^{2a_{1}})$ whenever $59\mid\sigma(5^{2a_{1}})$, which is absurd as $p_{i}\neq 35671$, for all $1\leq i\leq 6$. Therefore, no prime factor of $n$ can divides $\sigma(5^{2a_{1}})$ that is $\sigma(5^{2a_{1}})$ has prime factors other than $3$, $p_{1}=5$, $p_{2}$, …,$p_{5}$, $p_{6}$. This proves that $n$ can’t be a friend of $10$.
Subcase-9: For chain(9), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 19^{2a_{4}}\cdot 29^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $31\leq p_{6}\leq 37$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Now there are only three prime factors of $n$, $p_{3}=11$, $p_{4}=19$, $p_{6}=31$ that can divide $\sigma(5^{2a_{1}})$. By the same argument given in Subcase-8 we can conclude that $n$ can’t be a friend of $10$.
Subcase-10: For chain(10), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 17^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $19\leq p_{6}<\infty$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. But by Lemma 2.2, $I(n)>I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 17^{2a_{5}})\geq I(5^{2}\cdot 7^{2}\cdot 11^{2}\cdot 13^{2}\cdot 17^{2})=\frac{31}{25}\cdot\frac{57}{49}\cdot\frac{133}{121}\cdot\frac{183}{169}\cdot\frac{307}{289}>\frac{9}{5}$. Therefore $n$ can’t be a friend of $10$.
Subcase-11: For chain(11),$n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 19^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $23\leq p_{6}<\infty$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. But by Lemma 2.2 $I(n)>I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 19^{2a_{5}})\geq I(5^{2}\cdot 7^{2}\cdot 11^{2}\cdot 13^{2}\cdot 19^{2})=\frac{31}{25}\cdot\frac{57}{49}\cdot\frac{133}{121}\cdot\frac{183}{169}\cdot\frac{381}{361}>\frac{9}{5}$. Therefore $n$ can’t be a friend of $10$.
Subcase-12: For chain(12), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 23^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $29\leq p_{6}<\infty$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. If $22a_{1}>$ 2 then by Lemma 2.2, $I(n)>I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 23^{2a_{5}})\geq I(5^{4}\cdot 7^{2}\cdot 11^{2}\cdot 13^{2}\cdot 23^{2})=\frac{781}{625}\cdot\frac{57}{49}\cdot\frac{133}{121}\cdot\frac{183}{169}\cdot\frac{381}{361}>\frac{9}{5}$. Therefore $2a_{1}\leq 2$ that is $22a_{1}=$ 2. Since $\sigma(5^{2a_{1}})=\sigma(5^{2})=31$, $p_{6}$ must be $31$ but then $I(n)=I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 23^{2a_{5}}\cdot 31^{2a_{6}})\geq I(5^{2}\cdot 7^{2}\cdot 11^{2}\cdot 13^{2}\cdot 23^{2}\cdot 31^{2})=\frac{31}{25}\cdot\frac{57}{49}\cdot\frac{133}{121}\cdot\frac{183}{169}\cdot\frac{381}{361}\cdot\frac{993}{961}>\frac{9}{5}$. Therefore $n$ can’t be a friend of $10$.
We will discuss Subcase-13 at the very end, as it is a little bit more tricky than others. Before that, we will discuss the remaining subcases.
Subcase-14: For chain(14), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 31^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $37\leq p_{6}\leq 421$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Now $13\mid\sigma(n)$ if $p_{6}\in${53, 61, 79, 107, 113, 131, 139, 157, 191, 211, 263, 269, 313, 347, 367, 373, 419} that is for other values of $p_{6}$, $13\nmid\sigma(n)$. If $p_{6}\in$53,61,107,113,157,263,313,347,367,373}, then $p_{3}=11$ and $p_{5}=31$ only can divides $\sigma(5^{2a_{1}})$. But if $11\mid\sigma(5^{2a_{1}})$ then $71\mid\sigma(5^{2a_{1}})$ and if $31\mid\sigma(5^{2a_{1}})$ then $829\mid\sigma(5^{2a_{1}})$. Therefore $p_{3}=11$, $p_{5}=31$ can’t divide $\sigma(5^{2a_{1}})$ as for both the cases $p_{6}$ must be $71$ and $829$ respectively but $p_{6}$ neither be $71$ nor be $829$. Therefore $p_{6}$ must be one of $79,131,139,191,211,269,419$. If $p_{6}$ is one of them but does not divide $\sigma(5^{2a_{1}})$ then one of $11$, $31$ must divides $\sigma(5^{2a_{1}})$ but as previously discussed, it is not possible. Therefore, $p_{6}$ must divides $\sigma(5^{2a_{1}})$, but if $p_{6}=79\mid\sigma(5^{2a_{1}})$ then $305175781\mid\sigma(5^{2a_{1}})$, if $p_{6}=131$ or $211\mid\sigma(5^{2a_{1}})$ then $71\mid\sigma(5^{2a_{1}})$, if $p_{6}=139\mid\sigma(5^{2a_{1}})$ then $8971\mid\sigma(5^{2a_{1}})$, if $p_{6}=191\mid\sigma(5^{2a_{1}})$ then $6271\mid\sigma(5^{2a_{1}})$, if $p_{6}=269\mid\sigma(5^{2a_{1}})$ then $1609\mid\sigma(5^{2a_{1}})$, if $p_{6}=419\mid\sigma(5^{2a_{1}})$ then $191\mid\sigma(5^{2a_{1}})$. In each case, we are getting prime factors of $\sigma(5^{2a_{1}})$ other than $3$, $p_{1}=5$, $p_{2}$,…, $p_{6}$ and this is absurd by Remark 4.1. Therefore, no prime factor of $n$ can divides $\sigma(5^{2a_{1}})$ that is $\sigma(5^{2a_{1}})$ has prime factors other than $3$, $p_{1}=5$, $p_{2}$,…, $p_{5},p_{6}$. This proves that $n$ can’t be a friend of $10$.
Subcase-15: For chain(15), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 37^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $41\leq p_{6}\leq 127$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Now $13\mid\sigma(n)$ if $p_{6}\in$53,61,79,107,113} that is for other values of $p_{6}$, $13\nmid\sigma(n)$. If $p_{6}\in$53,61,107,113} we can use same argument given in Subcase-8 to show that, $n$ can’t be a friend of $10$. Therefore we assume $p_{6}$ is $79$. Now if $p_{6}$ does not divide $\sigma(5^{2a_{1}})$ then $11\mid\sigma(5^{2a_{1}})$ which allows $71\mid\sigma(5^{2a_{1}})$ but this is absurd by the fact that $p_{i}\neq 71$, for all $1\leq i\leq 6$. Therefore $p_{6}=79$ must divides $\sigma(5^{2a_{1}})$. But $31\mid\sigma(5^{2a_{1}})$ whenever $79\mid\sigma(5^{2a_{1}})$ and this shows that $p_{6}=79$ can’t divides $\sigma(5^{2a_{1}})$ as $p_{i}\neq 31$, for all $1\leq i\leq 6$. Therefore, no prime factor of $n$ can divides $\sigma(5^{2a_{1}})$ that is $\sigma(5^{2a_{1}})$ has prime factors other than $3$, $p_{1}=5$, $p_{2}$,…,$p_{5}$, $p_{6}$. This proves that $n$ can’t be a friend of $10$.
Subcase-16: For chain(16), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 41^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $43\leq p_{6}\leq 89$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Now $13\mid\sigma(n)$ if $p_{6}\in$53,61,79} that is for other values of $p_{6}$, $13\nmid\sigma(n)$. If $p_{6}\in$53,61} we can use same argument given in Subcase-8 to show that, $n$ can’t be a friend of $10$. Therefore assume that $p_{6}=79$. In this case also we can use similar argument given in Subcase-15 to show that, $n$ can’t be a friend of $10$.
Subcase-17: For chain(17), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 43^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $47\leq p_{6}\leq 83$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Now $13\mid\sigma(n)$ if $p_{6}\in$53,61,79} that is for other values of $p_{6}$, $13\nmid\sigma(n)$. If $p_{6}\in$53,61} we can use same argument given in Subcase-8 to show that, $n$ can’t be a friend of $10$. Therefore assume that $p_{6}=79$. In this case also we can use similar argument given in Subcase-15 to show that, $n$ can’t be a friend of $10$.
Subcase-18: For chain(18), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 47^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $51\leq p_{6}\leq 73$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. Now $13\mid\sigma(n)$ if $p_{6}=53$ or $p_{6}=61$ that is for other values of $p_{6}$, $13\nmid\sigma(n)$. For any $p_{6}$, only $11\mid\sigma(5^{2a_{1}})$ but it allows $71\mid\sigma(5^{2a_{1}})$ but this is absurd as, $p_{i}\neq 71$, for all $1\leq i\leq 6$. Therefore $n$ can’t be a friend of $10$.
Subcase-19: For chain(19), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 53^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $59\leq p_{6}\leq 61$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. If $p_{6}$ is other than 59 then only $p_{3}=11$ can divides $\sigma(5^{2a_{1}})$, but if it does then $71\mid\sigma(5^{2a_{1}})$, which is absurd. Therefore assume that $p_{6}=59$. Now $p_{3}=11$, $p_{6}=59$ can divides $\sigma(5^{2a_{1}})$ but using similar argument given in Subcase-8 we conclude that $n$ can’t be a friend of $10$.
Subcase-13: For chain(13), $n=5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}}\cdot 29^{2a_{5}}\cdot p_{6}^{2a_{6}}$ where $31\leq p_{6}\leq 20731$ and $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$, $a_{5}$, $a_{6}$ are positive integers. It is clear that $3$ must divides one of $\sigma(p_{i}^{2a_{i}})$ for $1\leq i\leq 6$ as $9\mid\sigma(n)$. But $3\nmid\sigma(p_{i}^{2a_{i}})$ for $i=1,3,5$. Therefore $3$ can divides $\sigma(7^{2a_{2}})$, $\sigma(13^{2a_{4}})$ and $\sigma(p_{6}^{2a_{6}})$. But if $3\mid\sigma(7^{2a_{2}})$ then $19\mid\sigma(7^{2a_{2}})$ but this is impossible as $p_{6}\geq 31$ and $p_{i}\neq 19$, for $1\leq i\leq 5$. Now if $3\mid\sigma(13^{2a_{4}})$ then $61\mid\sigma(13^{2a_{4}})$ that is $p_{6}$ must be $61$. Then $11$ is the only prime factor of $n$ which can divides $\sigma(5^{2a_{1}})$, but $11\mid\sigma(5^{2a_{1}})$ implies $71\mid\sigma(5^{2a_{1}})$ which is impossible. Hence $3$ must divides $\sigma(p_{6}^{2a_{6}})$. Again $5$ must divides one of $\sigma(p_{i}^{2a_{i}})$ for $1\leq i\leq 6$. But $5\nmid\sigma(p_{i}^{2a_{i}})$ for $i=1,2,4,5$. Therefore $5$ can divides $\sigma(11^{2a_{3}})$ and $\sigma(p_{6}^{2a_{6}})$. If $5\mid\sigma(11^{2a_{3}})$ then $3221\mid\sigma(11^{2a_{3}})$ which implies that $p_{6}=3221$ but this is absurd as $3\nmid\sigma(3221^{2a_{6}})$. Hence $5$ must divides $\sigma(p_{6}^{2a_{6}})$. Again $13$ must divides one of $\sigma(p_{i}^{2a_{i}})$ for $1\leq i\leq 6$, but $13\nmid\sigma(p_{i}^{2a_{i}})$ for $i=1,2,3,4$. Therefore $13$ can divides $\sigma(29^{2a_{4}})$ and $\sigma(p_{6}^{2a_{6}})$. If $13\mid\sigma(29^{2a_{5}})$ then $67\mid\sigma(29^{2a_{5}})$ that is $p_{6}$ must be $67$ but this is impossible as $5\nmid\sigma(67^{2a_{6}})$. Therefore we have $3$,$5$,$13$ divide $\sigma(p_{6}^{2a_{6}})$. Since $3\mid\sigma(p_{6}^{2a_{6}})$ and $5\mid\sigma(p_{6}^{2a_{6}})$ we have $p_{6}\equiv 1\pmod{30}$. Also we have $13\mid\sigma(p_{6}^{2a_{6}})$, and $p_{6}>13$ therefore by Theorem 1.5 we have $p_{6}\equiv 1\pmod{13}$ or $p_{6}\equiv 3\pmod{13}$ or $p_{6}\equiv 9\pmod{13}$ after solving all conditions on $p_{6}$ we obtain, $p_{6}\equiv 1\pmod{390}$ or $p_{6}\equiv 61\pmod{390}$ or $p_{6}\equiv 211\pmod{390}$. If $p_{6}\equiv 1\pmod{390}$ and since $31\leq p_{6}\leq 20731$, $p_{6}\in$1171, 1951, 2341, 2731, 3121, 3511, 5851, 7411, 8191, 8581, 8971, 10141, 10531, 11311, 11701, 14431, 14821, 15601, 15991, 16381, 17551, 19501, 19891}. TABLE 1 contains all such $p_{6}$ along with $p$ and $p^{*}$ such that whenever $p\mid\sigma(p_{6}^{2a_{6}})$ then $p^{*}\mid\sigma(p_{6}^{2a_{6}})$ also. If $p_{6}\equiv 61\pmod{390}$ and since $31\leq p_{6}\leq 20731$, $p_{6}\in$ 61, 1231, 1621, 2011, 2791, 3181, 3571, 5521, 6301, 6691, 8641, 9421, 9811, 12541, 13711, 15271, 15661, 16831, 20341, 20731 }. TABLE 2 contains all such $p_{6}$ along with $p$ and $p^{*}$ such that whenever $p\mid\sigma(p_{6}^{2a_{6}})$ then $p^{*}\mid\sigma(p_{6}^{2a_{6}})$ also. If $p_{6}\equiv 211\pmod{390}$ and since $31\leq p_{6}\leq 20731$, $p_{6}\in$ 211, 601, 991, 1381, 2161, 2551, 3331, 4111, 5281, 6451, 6841, 7621, 8011, 9181, 11131, 12301, 14251, 15031, 16981, 17761, 18541, 20101 }.TABLE 3 contains all such $p_{6}$ along with $p$ and $p^{*}$ such that whenever $p\mid\sigma(p_{6}^{2a_{6}})$ then $p^{*}\mid\sigma(p_{6}^{2a_{6}})$ also.
$p_{6}$ $p$ $p^{*}$
$1171$ $3$ $65353$
$1951$ $3$ $79$
$2341$ $3$ $37$
$2731$ $3$ $61$
$3121$ $5$ $521$
$3511$ $5$ $61$
$5851$ $5$ $181$
$7411$ $5$ $31$
$8191$ $3$ $22366891$
$8581$ $3$ $31$
$8971$ $3$ $271$
$10141$ $3$ $43$
$10531$ $3$ $157$
$11311$ $3$ $37$
$11701$ $3$ $97$
$14431$ $3$ $157$
$14821$ $5$ $271$
$15601$ $5$ $31$
$15991$ $5$ $61$
$16381$ $3$ $1237$
$17551$ $3$ $31$
$19501$ $3$ $163$
$19891$ $3$ $331$
Table 1.
$p_{6}$ $p$ $p^{*}$
$61$ $3$ $97$
$1231$ $3$ $37$
$1621$ $3$ $9631$
$2011$ $3$ $14821$
$2791$ $3$ $199807$
$3181$ $5$ $61$
$3571$ $5$ $101$
$5521$ $5$ $71$
$6301$ $5$ $31$
$6691$ $3$ $79$
$8641$ $5$ $3044081$
$9421$ $3$ $631$
$12541$ $5$ $151$
$13711$ $3$ $61$
$15271$ $3$ $43$
$15661$ $3$ $37$
$16831$ $3$ $109$
$20341$ $3$ $31$
$20731$ $5$ $251$
Table 2.
$p_{6}$ $p$ $p^{*}$
$211$ $3$ $37$
$601$ $3$ $9277$
$991$ $3$ $277$
$1381$ $5$ $811$
$2161$ $3$ $119797$
$2551$ $3$ $79$
$3331$ $5$ $41$
$4111$ $5$ $491$
$5281$ $3$ $715237$
$6451$ $3$ $152461$
$6841$ $5$ $61$
$7621$ $3$ $223$
$8011$ $5$ $41$
$9181$ $3$ $31$
$11131$ $5$ $31$
$12301$ $3$ $31$
$14251$ $3$ $5207827$
$15031$ $3$ $199$
$16981$ $3$ $7394137$
$17761$ $3$ $379$
$18541$ $3$ $79$
$20101$ $3$ $37$
Table 3.
This shows that for any $p_{6}$, we are getting some $p^{*}\not\in$3, $p_{1}=5$, $p_{2}$,…, $p_{6}$} such that $p^{*}\mid\sigma(p_{6}^{2a_{6}})$. In other words $\sigma(p_{6}^{2a_{6}})$ has other prime factors than $3$, $p_{1}=5$, $p_{2}$,…, $p_{6}$ which is absurd by Remark 4.1. Therefore $n$ can’t be a friend of $10$.
This proves that $n$ can’t have exactly $6$ distinct prime factors. Hence using Proposition 3.1, Proposition 3.2, Remark 4.2, Corollary 1.6 we prove the theorem.∎
5. Conclusion
We can of course use this method to check whether $n$ has seven or more distinct prime factors or not and we can predict the properties of prime factors of $n$ using the proven theorems. $10$ is not the only number whose classification is unknown to us. In fact the status of $14$, $15$, $20$ and many others are active topics for Research. It may be proving whether $10$ has a friend or not is as much as difficult as finding an odd perfect number. However computer search shows that if $10$ has a friend then its smallest friend will be at least $10^{30}$. The list of values of $f_{p}^{q}$ used in proving Theorem 1.7 can be found below.
$f_{31}^{5}=3$ $f_{11}^{5}=5$ $f_{71}^{5}=5$ $f_{19}^{5}=9$ $f_{829}^{5}=9$ $f_{305175781}^{5}=13$
$f_{191}^{5}=19$ $f_{6271}^{5}=19$ $f_{19}^{5}=9$ $f_{8971}^{5}=23$ $f_{59}^{5}=29$ $f_{35671}^{5}=29$
$f_{211}^{5}=35$ $f_{79}^{5}=39$ $f_{131}^{5}=65$ $f_{269}^{5}=67$ $f_{1609}^{5}=67$ $f_{139}^{5}=69$
$f_{419}^{5}=209$ $f_{3}^{7}=3$ $f_{19}^{7}=3$ $f_{9}^{7}=9$ $f_{37}^{7}=9$ $f_{5}^{11}=5$
$f_{3221}^{11}=5$ $f_{3}^{13}=3$ $f_{61}^{13}=3$ $f_{3}^{19}=3$ $f_{127}^{19}=3$ $f_{13}^{29}=3$
$f_{67}^{29}=3$ $f_{3}^{61}=3$ $f_{97}^{61}=3$ $f_{3}^{211}=3$ $f_{37}^{211}=3$ $f_{3}^{601}=3$
$f_{9277}^{601}=3$ $f_{3}^{991}=3$ $f_{277}^{991}=3$ $f_{5}^{1021}=5$ $f_{41}^{1021}=5$ $f_{1451}^{1021}=5$
$f_{3}^{1171}=3$ $f_{65353}^{1171}=3$ $f_{3}^{1231}=3$ $f_{37}^{1231}=3$ $f_{5}^{1381}=5$ $f_{811}^{1381}=5$
$f_{5}^{1531}=5$ $f_{691}^{1531}=5$ $f_{3}^{1621}=3$ $f_{9631}^{1621}=3$ $f_{3}^{1951}=3$ $f_{79}^{1951}=3$
$f_{3}^{2011}=3$ $f_{14821}^{2011}=3$ $f_{3}^{2161}=3$ $f_{119797}^{2161}=3$ $f_{3}^{2341}=3$ $f_{37}^{2341}=3$
$f_{3}^{2551}=3$ $f_{79}^{2551}=3$ $f_{3}^{2731}=3$ $f_{61}^{2731}=3$ $f_{3}^{2791}=3$ $f_{199807}^{2791}=3$
$f_{5}^{3121}=5$ $f_{521}^{3121}=5$ $f_{5}^{3181}=5$ $f_{61}^{3181}=5$ $f_{5}^{3331}=5$ $f_{41}^{3331}=5$
$f_{5}^{3511}=5$ $f_{61}^{3511}=5$ $f_{5}^{3571}=5$ $f_{101}^{3571}=5$ $f_{5}^{4111}=5$ $f_{491}^{4111}=5$
$f_{3}^{5281}=3$ $f_{715237}^{5281}=3$ $f_{5}^{5521}=5$ $f_{71}^{5521}=5$ $f_{5}^{5851}=5$ $f_{181}^{5851}=5$
$f_{5}^{6301}=5$ $f_{31}^{6301}=5$ $f_{3}^{6451}=3$ $f_{152461}^{6451}=3$ $f_{3}^{6691}=3$ $f_{79}^{6691}=3$
$f_{5}^{6841}=5$ $f_{61}^{6841}=5$ $f_{5}^{7411}=5$ $f_{31}^{7411}=5$ $f_{3}^{7621}=3$ $f_{223}^{7621}=3$
$f_{5}^{8011}=5$ $f_{41}^{8011}=5$ $f_{3}^{8191}=3$ $f_{22366891}^{8191}=3$ $f_{3}^{8581}=3$ $f_{31}^{8581}=3$
$f_{5}^{8641}=5$ $f_{3044081}^{8641}=5$ $f_{3}^{8971}=3$ $f_{271}^{8971}=3$ $f_{3}^{9181}=3$ $f_{31}^{9181}=3$
$f_{3}^{9421}=3$ $f_{631}^{9421}=3$ $f_{3}^{10141}=3$ $f_{43}^{10141}=3$ $f_{3}^{10531}=3$ $f_{157}^{10531}=3$
$f_{5}^{11131}=5$ $f_{31}^{11131}=5$ $f_{3}^{11311}=3$ $f_{37}^{11311}=3$ $f_{3}^{11701}=3$ $f_{97}^{11701}=3$
$f_{3}^{12301}=3$ $f_{31}^{12301}=3$ $f_{5}^{12541}=5$ $f_{151}^{12541}=5$ $f_{3}^{13711}=3$ $f_{61}^{13711}=3$
$f_{3}^{14251}=3$ $f_{5207827}^{14251}=3$ $f_{3}^{14431}=3$ $f_{157}^{14431}=3$ $f_{5}^{14821}=5$ $f_{271}^{14821}=5$
$f_{3}^{15031}=3$ $f_{199}^{15031}=3$ $f_{3}^{15271}=3$ $f_{43}^{15271}=3$ $f_{5}^{15601}=5$ $f_{127}^{15601}=5$
$f_{3}^{15661}=3$ $f_{37}^{15661}=3$ $f_{5}^{15991}=5$ $f_{61}^{15991}=5$ $f_{3}^{16381}=3$ $f_{1237}^{16381}=3$
$f_{3}^{16831}=3$ $f_{109}^{16831}=3$ $f_{3}^{16981}=3$ $f_{7394137}^{16981}=3$ $f_{3}^{17551}=3$ $f_{31}^{17551}=3$
$f_{3}^{17761}=3$ $f_{379}^{17761}=3$ $f_{3}^{18541}=3$ $f_{79}^{18541}=3$ $f_{3}^{19501}=3$ $f_{163}^{19501}=3$
$f_{3}^{19891}=3$ $f_{331}^{19891}=3$ $f_{3}^{20101}=3$ $f_{37}^{20101}=3$ $f_{3}^{20341}=3$ $f_{31}^{20341}=3$
$f_{5}^{20731}=5$ $f_{251}^{20731}=5$
Table 4. List of $f_{p}^{q}$